Find the area of the rectangles whose sides are:
 3 cm and 4 cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that
Area of rectangle = Length × Breadth
l = 3 cm and b = 4 cm
Area = l × b = 3 × 4
= 12 cm^2

Thanks